OpenStudy (anonymous):

identify the slope and y-intercept of 3x-2y=7 My answer: the slope is 3/2 and the y intercept is 7/2. Is this correct?

jagr2713 (jagr2713):

Y intercept is close but you forgot your negative sign

OpenStudy (anonymous):

so slope is 3/2 and y intercept is -7/2?

jagr2713 (jagr2713):

Correct!

OpenStudy (anonymous):

okay thanks

jagr2713 (jagr2713):

yw

OpenStudy (anonymous):

3x - 2y = 7; subtract 3x from each side 3x - 3x - 2y = 7 - 3x; simplify and rearrange -2y = -3x + 7; divide each term by -2 -2y/-2 = -3x/-2 + 7/-2; simplify wud u say this work is correct